# From Continual Learning to SGD and Back: Better Rates for Continual Linear Models

### v2

@@ 3.2 New Reduction: Kaczmarz Method and Stepwise-Optimal Stochastic Gradient Descent
Rank-$1$ data.
It is known that when each task contains just one row, each update in the Kaczmarz method corresponds to a gradient step on with a specific “normalizing” step size [51]. That is, since in rank-$1$ we have $\mathcal{L}_{\tau(t)}(\mathbf{w})=\frac{1}{2}\big{\|}{\mathbf{x}_{\tau({t})}^{\top}\mathbf{w}-y_{\tau({t})}}\big{\|}^{2}$, the Kaczmarz updates hold
@@
Reduction 2 (Block Kaczmarz $\Rightarrow$ SGD).
In the realizable case (Assumption 2.2) under any ordering $\tau$, the block Kaczmarz method is equivalent to SGD with a step size of $\eta=1$, applied w.r.t. a convex, $1$-smooth least squares objective: $\big{\{}f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\left({\mathbf{X}_{m}\mathbf{w}-\mathbf{y}_{m}}\right)}\right\|^{2}\big{\}}_{m=1}^{T}$. That is, the iterates $\mathbf{w}_{0},\dots,\mathbf{w}_{k}$ of Schemes 2 and 3 coincide.
The reduction follows from the next lemma, which states the convexity and smoothness of $f_{m}$ and expresses the gradient $\nabla_{\mathbf{w}}f_{\tau(t)}$, subsequently substituted into $\big{(}{\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})}\big{)}$ to complete the proof. The proof for the lemma is given in App. B.
@@
Consider any realizable task collection such that ${\mathbf{X}_{m}\mathbf{w}_{\star}=\mathbf{y}_{m}},\forall m\in\left[{T}\right]$. Define ${f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\left({\mathbf{X}_{m}\mathbf{w}-\mathbf{y}_{m}}\right)}\right\|^{2}}$. Then, $\forall m\in\left[{T}\right],\mathbf{w}\in\mathbb{R}^{d}$
@@
    Upper bound: $\mathcal{L}_{m}(\mathbf{w})\leq R^{2}f_{m}(\mathbf{w})\triangleq\max_{m^{\prime}\in\left[{T}\right]}\left\|{\mathbf{X}_{m^{\prime}}}\right\|^{2}f_{m}(\mathbf{w})$
@@
    Gradient:       $\nabla_{\mathbf{w}}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\mathbf{w}-\mathbf{X}_{m}^{+}\mathbf{y}_{m}$
@@ 4 Rates for Random-Order Continual Linear Regression and Kaczmarz
This section focuses on improving the best upper bound known in prior continual learning literature on random orderings, summarized in Table 1. Specifically, for with-replacement random orderings, Evron et al. [15] proved a forgetting rate of $\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]=\mathcal{O}\!\left({\frac{d-\bar{r}}{k}}\right)$ where $\bar{r}\triangleq\frac{1}{T}\sum_{m}\operatorname{rank}(\mathbf{X}_{m})$. Notably, this rate depends on the problem dimensionality $d$, raising concerns when generalizing insights from linear models to deep networks, which are often heavily overparameterized (e.g., in the NTK regime). Encouragingly, that paper only provided (implicitly) a $1/k$ lower bound for the worst-case forgetting, calling for further research into narrowing this gap.
We tighten the existing rate’s problem-dependent term from $\big{(}{d-\bar{r}}\big{)}$ to $\min\big{(}{\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\big{)}$ and also prove a problem-independent rate of ${1/\sqrt[4]{k}}$. Finally, we provide the first rates for without-replacement orderings, isolating the effect of randomness versus repetition.
Table 1: Forgetting and Loss Rates in Continual Linear Regression (and Block Kaczmarz). Upper bounds apply to any $T$ realizable tasks (or blocks). Lower bounds indicate worst cases, i.e., specific constructions. Random ordering bounds apply to the expected forgetting (or loss). We omit mild constant multiplicative factors and an unavoidable $\left\|{\mathbf{w}_{\star}}\right\|^{2\!}R^{2}$ term. Finally, ${a\wedge b\triangleq\min(a,b)}$. Recall: $k=\,$iterations; $d=\,$dimensionality; $\bar{r},r_{\max}=\,$average and maximum data matrix ranks.
@@
Evron et al. [15] Upper $\frac{d-\bar{r}}{k}$  $\frac{T^{2}}{\sqrt{k}}\wedge\frac{T^{2}(d-r_{\max})}{k}$
Kong et al. [34] Upper $\frac{T^{3}}{k}$
Ours (2025) Upper $\frac{1}{\sqrt[4]{k}}\wedge\frac{\sqrt{d-{\bar{r}}}}{k}\wedge\frac{\sqrt{T\bar{r}}}{k}$ $\frac{1}{\sqrt[4]{T}}\wedge\frac{d-{\bar{r}}}{T}$
Evron et al. [15] Lower $\frac{1}{k}$ (*) $\frac{1}{T}$ (*) $\frac{T^{2}}{k}$
(*) They did not explicitly provide such lower bounds, but the $T\!=\!2$ tasks construction from their proof of Theorem 10, can yield a $\Theta({1}/{k})$ random behavior by cloning those $2$ tasks $\left\lfloor{T/2}\right\rfloor$ times for any general $T$.
@@ 4.1 A Parameter-Dependent $\mathcal{O}(1/k)$ Rate
The $\left\|{\mathbf{w}_{\star}}\right\|^{2}\!R^{2}$ scaling term is generally unavoidable. Our proof, given in App. C, is related to a recent work [23] that characterized the weak error (resembling our loss) by analyzing a linear map. Unlike ours, the polynomial rates they derive involve matrix properties related to the condition number.
@@ 4.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
Next, we present a forgetting rate independent on the dimensionality, rank, and number of tasks. This is crucial in highly overparameterized regimes which are connected to deep neural networks.
@@
[…]
We prove this result in Section D.1 by leveraging the connections between CL and SGD. Specifically, in Section 3 we showed that continual linear regression is equivalent to SGD with a step size of exactly $1$on a related least squares objective that bounds the original continual learning loss. Our result then follows from our novel last-iterate SGD bounds that, crucially, apply even to that specific step size. To ease readability, here we focused on a CL perspective, deferring last-iterate analysis to Section 5.
@@ 4.3 Random Task Orderings Without Replacement
Evron et al. [15] suggested defining forgetting as “catastrophic” only when $\lim_{k\to\infty}\mathbb{E}\left[F_{\tau}(k)\right]>0$.They presented such an adversarial case with a deterministic task ordering where $k=T\to\infty$, and showed that task recurrence, under cyclic or random orderings, mitigates forgetting. So far, in random orderings, it was hard to isolate the effect of randomness from that of repetitions. It was thus unclear whether catastrophic forgetting can be alleviated by randomly permuting the tasks. Below, we provide the first result demonstrating that no recurrence is needed under random orderings.
@@
The proof of the dimensionality-dependent term is similar to the one of the with-replacement case, given in Section D.1.2 of Evron et al. [15], but requires a more careful upper bound on the (in-sample) forgetting. The proof of the dimensionality-independent term again relies on last-iterate analysis, as presented in Section E.2. Both proofs are given in Section D.2.
App. A discusses connections between our result above and related areas like shuffle SGD.
@@ 5 Last-Iterate SGD Bounds for Linear Regression
In this self-contained section, we derive last-iterate guarantees for SGD in the realizable stochastic least squares setup. Motivated by the connection with continual regression discussed in Section 3, we focus on regression problems that are $\beta$-smooth individually, and obtain upper bounds for the last SGD iterate that apply for a significantly wider range of step sizes compared to prior art [67]. Notably, this is the first time convergence of SGD in this setup is established for a range of step sizes completely independent of the optimization horizon. Table 3 in App. A compares our bounds with related work and classical results in the field.
Recent work has analyzed SGD in the realizable least squares setting, with most assuming a somewhat more general noise model [18, 68, 4, 73, 67, 69]. These studies are primarily motivated by connections to deep networks in the overparameterized regime [45], where models are expressive enough to perfectly fit the training data. With the exception of Varre et al. [67], most of these works focus on non-fixed step sizes and/or provide guarantees for the average iterate. See App. A for further discussion. Here, we study a stochastic, jointly realizable least squares problem, as defined next.
Setup 1.
@@
where $\mathbf{A}_{i}\in\mathbb{R}^{n_{i}\times d},\mathbf{b}_{i}\in\mathbb{R}^{n_{i}},~{}\forall i\in\mathcal{I}$. We specifically focus on $\beta$-smooth functions, i.e., ${\left\|{\mathbf{A}_{i}^{\top}\mathbf{A}_{i}}\right\|\leq\beta},\forall{i\in\mathcal{I}}$, under a realizable assumption, i.e., $\exists\mathbf{w}_{\star}\in\mathbb{R}^{d}:\bar{f}(\mathbf{w}_{\star})=0$.
@@
The important feature of Theorem 5.1 is the $(2-\eta\beta)$ factor in the denominator, replacing the common $(1-\eta\beta)$ of the standard analysis. This difference makes our theorem applicable to the continual regression setting which requires setting $\eta=1/\beta$ (Reduction 2). In addition, for $\eta={1}/{\left({\beta\log T}\right)}$, we recover the near-optimal rate obtained by Varre et al. [67], i.e., $\mathbb{E}\bar{f}(\mathbf{w}_{T})=\mathcal{O}\left({\frac{\beta D^{2}\log T}{T}}\right).$
@@
In Section E.2, we extend Theorem 5.1 to the setting of SGD without replacement. The proof leverages algorithmic stability for SGD [6, 61, 25], focusing on a variant tailored to without-replacement sampling [63, 35]. In particular, we establish a new bound for this variant in the smooth and realizable regime, which has not appeared in prior work.
@@
The second and main step of the analysis is to relate the loss of the last SGD iterate to the regret of the algorithm. For this, we carefully adapt an existing approach for last-iterate convergence in the non-smooth case [62]. The result, given below, is slightly more general to accommodate without-replacement sampling, addressed in the next section.
@@ 6.1 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for General Projections Onto Convex Sets
Projections Onto Convex Sets (POCS) is a classical method that iteratively projects onto closed convex sets to find a point in their intersection [22, 7]. Formally,
@@
Reduction 3 (POCS $\Rightarrow$ SGD).
@@
To the best of our knowledge, this is the first universal rate in the POCS literature, independent of problem parameters such as its regularity or complexity, as demonstrated next in Section 6.2. Universal rates are only achievable when analyzing individual distances, i.e., $f_{m}(w)=\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{m})=\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})}\right\|^{2}$, rather than the distance to the intersection, i.e., $\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{\star})$. In machine learning, the squared distance from individual sets is linked to important losses like MSE in regression or squared hinge loss in classification [15, 16], naturally leading to our next continual model.

### v3

@@ 3.2 New Reduction: Kaczmarz Method and Stepwise-Optimal Stochastic Gradient Descent
Rank-11 data.
It is known that when each task contains just one row, each update in the Kaczmarz method corresponds to a gradient step on with a specific “normalizing” step size (Needell et al., 2014). That is, since in rank-$1$ we have $\mathcal{L}_{\tau(t)}(\mathbf{w})=\frac{1}{2}\big\|{\mathbf{x}_{\tau({t})}^{\top}\mathbf{w}-y_{\tau({t})}}\big\|^{2}$, Kaczmarz updates hold
@@
Reduction 2 (Block Kaczmarz $\Rightarrow$ SGD)
In the realizable case (Section 2) under any ordering $\tau$, the block Kaczmarz method is equivalent to SGD with a step size of $\eta=1$, applied w.r.t. a convex, $1$-smooth least squares objective: $\big\{f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\left({\mathbf{X}_{m}\mathbf{w}-\mathbf{y}_{m}}\right)}\right\|^{2}\big\}_{m=1}^{T}$. That is, the iterates $\mathbf{w}_{0},\dots,\mathbf{w}_{k}$ of Schemes 2 and 3 coincide.
Intuitively, the $\mathbf{X}_{m}^{+}$ term in the modified objectives $\{f_{m}\}$ generalizes the normalizing step size from the rank-1 case, fitting all directions in the current block precisely with the same step size $\eta=1$.
The reduction above is key to our analysis flow (Figure 1) as it reveals that continual linear regression can be analyzed directly via SGD analysis. It follows from substituting the gradient from the next lemma into $({\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})})$ in Scheme 3. The lemma is proved in App. B.
@@
[…]
@@
    Upper bound: $\mathcal{L}_{m}(\mathbf{w})\leq R^{2}f_{m}(\mathbf{w})\triangleq\max_{m^{\prime}\in\left[{T}\right]}\left\|{\mathbf{X}_{m^{\prime}}}\right\|^{2}f_{m}(\mathbf{w})$ .
@@
    Gradient:    $\nabla_{\mathbf{w}}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\mathbf{w}-\mathbf{X}_{m}^{+}\mathbf{y}_{m}$ .
@@ 4 Rates for Random-Order Continual Linear Regression and Kaczmarz
This section improves the best known upper bound: for random orderings with replacement, Evron et al. (2022) proved a forgetting rate of $\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]=\mathcal{O}\!\left({\frac{d-\bar{r}}{k}}\right)$ where $\bar{r}\triangleq\frac{1}{T}\sum_{m}\operatorname{rank}(\mathbf{X}_{m})$. Notably, their rate depends on the dimensionality $d$, challenging the transfer of insights from linear models to highly overparameterized deep networks (e.g., via the NTK regime). Encouragingly, they only provided a worst-case lower bound of $1/k$, calling for further research to narrow this gap.
We tighten the existing problem-dependent rate from $\big({d-\bar{r}}\big)$ to $\min\big({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\big)$, and prove a problem-independent rate of $1/\sqrt[4]{k}$}. Finally, we provide the first rates for without-replacement orderings, isolating the effect of randomness versus repetition. See summary in the table below.
Table 1: Forgetting and Loss Rates in Continual Linear Regression (and Block Kaczmarz). Upper bounds apply to any $T$ realizable tasks (or blocks). Lower bounds indicate worst cases, i.e., specific constructions. Random ordering bounds apply to the expected forgetting (or loss). We omit mild constant multiplicative factors and an unavoidable $\left\|{\mathbf{w}_{\star}}\right\|^{2\!}R^{2}$ term. Finally, $a\wedge b\triangleq\min(a,b)$}. Recall: k=k=\,iterations; d=d=\,dimensionality; $\bar{r},r_{\max}=\,$average and maximum data matrix ranks.
@@
  Evron et al. (2022) Upper $\frac{d-\bar{r}}{k}$  $\frac{T^{2}}{\sqrt{k}}\wedge\frac{T^{2}(d-r_{\max})}{k}$
Kong et al. (2023) Upper $\frac{T^{3}}{k}$
Ours Upper $\frac{1}{\sqrt[4]{k}}\wedge\frac{\sqrt{d-{\bar{r}}}}{k}\wedge\frac{\sqrt{T\bar{r}}}{k}$ $\frac{1}{\sqrt[4]{T}}\wedge\frac{d-{\bar{r}}}{T}$
  Evron et al. (2022) Lower $\frac{1}{k}$ (*) $\frac{1}{T}$ (*) $\frac{T^{2}}{k}$
(*) They did not explicitly provide such lower bounds, but the 22-task construction from their proof of Theorem 10, can yield a $\Theta({1}/{k})$ random behavior by cloning those 22 tasks $\left\lfloor{T/2}\right\rfloor$ times for any general $T$.
@@ 4.1 A Parameter-Dependent $\mathcal{O}(1/k)$ Rate
Our proof, given in App. C, is related to a recent work (Guo et al., 2022) that characterizes the weak error (similar to our loss) by analyzing a linear map. Unlike ours, the polynomial rates they derive involve matrix properties related to the condition number.
@@
[…]
Remark 4.2 (The $\left\|{\mathbf{w}_{\star}}\right\|^{2}\!R^{2}$ Scaling Term).
All the rates we derive contain a multiplicative factor of $\|\mathbf{w}_{\star}\|^{2}R^{2}$, a generally unavoidable scaling term in linear regression. Prior work on continual learning has either normalized it away implicitly—e.g., by assuming $\left\|{\mathbf{w}_{\star}}\right\|^{2},R\leq 1$ (Evron et al., 2022)—or included it explicitly, as we do (Evron et al., 2023; Lin et al., 2023). The rate in Theorem 4.1 involves additional problem parameters, i.e., $T$, $d$, and $\bar{r}$, whereas the rate in Theorem 4.3 below is “universal” in the sense that it does not depend on any such parameter.
@@ 4.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
Next, we present a forgetting rate independent on the dimensionality, rank, and number of tasks. This independence is crucial in highly overparameterized regimes, as encountered in deep neural networks.
@@
$\mathbb{E}_{\tau}\!\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $\leq\frac{2\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{\sqrt[4]{k}}\,,\quad\quad\mathbb{E}_{\tau}\!\left[F_{\tau}(k)\right]\leq\frac{5\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{\sqrt[4]{k-1}}\,.$
We prove this result in Section D.1 by leveraging the connections between CL and SGD. Specifically, Section 3 showed that continual linear regression is equivalent to SGD with step size exactly $1$ on a related least squares objective that bounds the original continual learning loss. Our result then follows from our novel last-iterate SGD bounds that, crucially, apply even to this specific step size. To ease readability, we keep a CL perspective here and defer last-iterate analysis to Section 5.
@@ 4.3 Random Task Orderings Without Replacement
Evron et al. (2022) suggested that forgetting is ‘catastrophic’ only when $\lim_{k\to\infty}\mathbb{E}\left[F_{\tau}(k)\right]>0$, and presented such an adversarial case with a deterministic task ordering where $k=T\to\infty$. In contrast, they showed that cyclic or random task orderings mitigate forgetting, perhaps due to task repetition. So far, under random orderings, it has been difficult to disentangle the effect of randomness from that of repetition—i.e., whether their remedying impact arises from random permutation or repeated exposure. Below, we provide the first result demonstrating that randomly permuting tasks is sufficient to alleviate catastrophic forgetting.
@@
The proof is given in Section D.2. The dimensionality-dependent term parallels the with-replacement case in App. D.1.2 of Evron et al. (2022), but requires a refined upper bound on in-sample forgetting. The dimensionality-independent term again relies on last-iterate analysis, as presented in Section E.2.
In App. A, we discuss connections between our result above and areas like shuffle SGD.
@@ 5 Last-Iterate SGD Bounds for Linear Regression
In this self-contained section, we derive last-iterate guarantees for SGD in the realizable stochastic least squares setup. Motivated by the connection with continual regression discussed in Section 3, we focus on regression problems that are $\beta$-smooth individually, and obtain upper bounds for the last SGD iterate that apply for a significantly wider range of step sizes compared to prior art (Varre et al., 2021). Notably, this is the first time convergence of SGD in this setup is established for a range of step sizes completely independent of the optimization horizon. Table 3 in App. A compares our bounds with related work and classical results in the field.
Recent work has analyzed SGD in realizable (possibly noisy) least squares settings (Ge et al., 2019; Vaswani et al., 2019; Berthier et al., 2020; Zou et al., 2021; Varre et al., 2021; Wu et al., 2022). Realizable settings are primarily motivated by connections to deep networks in the overparameterized regime (Ma et al., 2018), where models are expressive enough to perfectly fit the training data. With the exception of Varre et al. (2021), most of these works focus on non-fixed step sizes and/or provide guarantees for the average iterate (see App. A for discussion). Similarly, here we study the following stochastic, jointly realizable least squares problem.
Setup 1
@@
where $\mathbf{A}_{i}\in\mathbb{R}^{n_{i}\times d},\mathbf{b}_{i}\in\mathbb{R}^{n_{i}},~\forall i\in\mathcal{I}$. We specifically focus on $\beta$-smooth functions, that is, ${\left\|{\mathbf{A}_{i}^{\top}\mathbf{A}_{i}}\right\|\leq\beta},\forall{i\in\mathcal{I}}$, under a realizable assumption, i.e., $\exists\mathbf{w}_{\star}\in\mathbb{R}^{d}:\bar{f}(\mathbf{w}_{\star})=0$.
@@
An important part of Theorem 5.1 is the $(2-\eta\beta)$ factor in the denominator, replacing a $(1-\eta\beta)$common in standard analysis. This difference makes our theorem applicable to the continual regression setting which requires setting $\eta=1/\beta$ (Reduction 2). In addition, for $\eta={1}/{\left({\beta\log T}\right)}$, we recover the near-optimal rate obtained by Varre et al. (2021), i.e., $\mathbb{E}\bar{f}(\mathbf{w}_{T})=\mathcal{O}\left({\frac{\beta D^{2}\log T}{T}}\right).$
@@
In Section E.2, we extend Theorem 5.1 to SGD without replacement. The proof leverages algorithmic stability for SGD (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010; Hardt et al., 2016), focusing on a variant tailored to without-replacement sampling (Sherman et al., 2021; Koren et al., 2022). In particular, we establish a new bound for this variant in the smooth and realizable regime, which has not appeared in prior work.
@@
The second and main step of the analysis is to relate the loss of the last SGD iterate to the regret of the algorithm. For this, we carefully adapt an existing approach for last-iterate convergence in the non-smooth case (Shamir and Zhang, 2013). The result, given below, is slightly more general to accommodate without-replacement sampling, addressed in the next section.
@@ 6.1 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for General Projections Onto Convex Sets
Projections Onto Convex Sets (POCS) is a classical method that iteratively projects onto closed convex sets to find a point in their intersection (Gubin et al., 1967; Boyd et al., 2003). Formally,
@@
Reduction 3 (POCS $\Rightarrow$ SGD)
@@
To the best of our knowledge, this is the first universal rate in the POCS literature, independent of problem parameters such as regularity or complexity, as demonstrated in Section 6.2. Universal rates are only achievable when analyzing individual distances, i.e., $f_{m}(w)=\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{m})=\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})}\right\|^{2}$, rather than the distance to the intersection, i.e., $\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{\star})$. In machine learning, squared distances from individual sets relate not only to MSE, but also to losses such as the squared hinge loss for classification (Evron et al., 2023), naturally leading to our next continual model.
Figure 2: Analysis Flow Leading to Our Improved Classification Rates.